### v1

# Skew braces with no proper left ideals

Abstract.
A skew brace $A=(A,\cdot,\circ)$ is said to be left-simple if $A\neq 1$ and it has no left ideal other than $1$ and $A$. The purpose of this paper is to give a partial classification of the finite left-simple skew braces. A result of Stefanello and Trappeniers implies that finite left-simple skew braces correspond precisely to minimal Hopf–Galois structures on finite Galois extensions of fields.
Key words and phrases:
skew brace, left ideal, minimal Hopf–Galois structure
2020 Mathematics Subject Classification:
Primary 20N99 16Y99, Secondary 12F10
1. Introduction
A skew brace is a set $A=(A,\cdot,\circ)$ equipped with two group operations $\cdot$ and $\circ$ such that the so-called brace relation
$a\circ(b\cdot c)=(a\circ b)\cdot a^{-1}\cdot(a\circ c)$
holds for all $a,b,c\in A$. For any $a\in A$, we write $a^{-1}$ and $\overline{a}$, respectively, for the inverses of $a$ in the groups $(A,\cdot)$ and $(A,\circ)$. We easily check that $(A,\cdot)$ and $(A,\circ)$ share the same identity $1$. For each $a\in A$, it is straightforward to verify that the map
$\lambda_{a}:A\longrightarrow A;\quad\lambda_{a}(x)=a^{-1}\cdot(a\circ x)$
is an automorphism of $(A,\cdot)$, and it is also not hard to show that
$\lambda:(A,\circ)\longrightarrow\mathrm{Aut}(A,\cdot);\quad a\mapsto\lambda_{a}$ (1.1)
is a group homomorphism. A left ideal of $A$ is a subgroup $I$ of $(A,\cdot)$ that is invariant under the action of $\lambda$, namely $\lambda_{a}(I)\subseteq I$ for all $a\in A$. Note that $I$ is a subgroup of $(A,\circ)$ in this case. We introduce a new definition.
Definition 1.1.
We shall say that a skew brace $A=(A,\cdot,\circ)$ is left-simple if $A\neq 1$ and it has no left ideal other than $1$ and $A$.
Analogous to the Classification of Finite Simple Groups, it is natural to ask for a classification of the finite left-simple skew braces. The purpose of this paper is to give a partial classification.
Let $A=(A,\cdot,\circ)$ be any finite left-simple skew brace. The additive group of $A$, namely $(A,\cdot)$, must be characteristically simple because its characteristic subgroups are all left ideals of $A$. Hence, we have $(A,\cdot)\simeq T^{n}$ for some finite simple group $T$ and $n\in\mathbb{N}$. We are able to prove a complete classification when $T$ is abelian and when $n=1$. We give a few fairly restrictive conditions that $A$ must satisfy when $T$ is non-abelian and $n\geq 2$.
Theorem 1.2.
Let $A=(A,\cdot,\circ)$ be a finite skew brace such that $(A,\cdot)\simeq\mathbb{F}_{p}^{n}$ for some prime $p$ and $n\in\mathbb{N}$. Then $A$ is left-simple if and only if $n=1$.
Proof.
We may regard the group homomorphism
$\lambda:(A,\circ)\longrightarrow\mathrm{Aut}(A,\cdot)\simeq\mathrm{GL}_{n}(\mathbb{F}_{p})$
as a linear representation of the finite $p$-group $(A,\circ)$ over the field $\mathbb{F}_{p}$ that has characteristic $p$. The left ideals of the skew brace $A$ then coincide with the $\lambda$-invariant subspaces of the vector space $(A,\cdot)\simeq\mathbb{F}_{p}^{n}$. This means that $A$ is left-simple if and only if $\lambda$ is irreducible. But it is well-known in modular representation theory (e.g. see [11, Proposition 6.2.1]) that $\lambda$ is irreducible if and only if it is the trivial representation. This proves the theorem. ∎
The following corollary is immediate from Theorem 1.2.
Corollary 1.3.
Up to isomorphism, the finite left-simple skew braces with abelian additive group are exactly $(\mathbb{F}_{p},+,+)$, where $p$ is any prime and ++ is the usual addition.
Theorem 1.4.
Let $A=(A,\cdot,\circ)$ be a finite skew brace such that $(A,\cdot)\simeq T^{n}$ for some non-abelian simple group $T$ and $n\in\mathbb{N}$. It is well-known that
$\mathrm{Aut}(T^{n})=\mathrm{Aut}(T)\wr S_{n}=\mathrm{Aut}(T)^{n}\rtimes S_{n},$
where $S_{n}$ denotes the symmetric group on $n$ letters.
  1. $(a)$
    In the case $n=1$, we have $A$ is left-simple if and only if $A$ is an almost trivial skew brace, namely, $a\circ b=b\cdot a$ for all $a,b\in A$.
  2. $(b)$
    In the case $n\geq 2$, if $A$ is left-simple, then $(A,\circ)$ is isomorphic to some subgroup of $\mathrm{Aut}(T^{n})$ whose projection onto $S_{n}$ is a transitive subgroup, and $\mathrm{Im}(\lambda)$ intersects trivially with $\mathrm{Inn}(A,\cdot)$.
Remark 1.5.
The motivation of this paper came from the study of minimal Hopf–Galois structures on finite Galois extensions. In the language of Hopf–Galois theory, our results above yield a partial classification of the minimal Hopf–Galois structures on finite Galois extensions (see Section 4).
2. Preliminaries
Let $A=(A,\cdot,\circ)$ be a skew brace. For each $a\in A$, the maps
$\lambda_{a}:A\longrightarrow A;$ $\quad\lambda_{a}(x)=a^{-1}\cdot(a\circ x)$
$\rho_{a}:A\longrightarrow A;$ $\quad\rho_{a}(x)=(a\circ x)\cdot a^{-1}$
are automorphisms of $(A,\cdot)$. Also, let
$\mathrm{conj}(a):A\longrightarrow A;\quad\mathrm{conj}(a)(x)=a\cdot x\cdot a^{-1}$
denote the inner automorphism of $(A,\cdot)$ induced by $a$, so that
$\rho_{a}=\mathrm{conj}(a)\lambda_{a},\quad\lambda_{a}=\mathrm{conj}(a^{-1})\rho_{a},\quad\mathrm{conj}(a)=\rho_{a}\lambda_{a}^{-1}$ (2.1)
hold. It is well-known that
$\lambda:(A,\circ)\longrightarrow\mathrm{Aut}(A,\cdot);$ $\quad a\mapsto\lambda_{a}$
$\rho:(A,\circ)\longrightarrow\mathrm{Aut}(A,\cdot);$ $\quad a\mapsto\rho_{a}$
are group homomorphisms. Also let
$\mathrm{conj}:(A,\cdot)\longrightarrow\mathrm{Aut}(A,\cdot);\quad a\mapsto\mathrm{conj}(a)$
denote the natural group homomorphism.
For the homomorphisms $\lambda$ and $\rho$, we consider the product
$\Gamma:=\mathrm{Im}(\lambda)\mathrm{Im}(\rho)$
of their images. The next proposition is essentially [10, Proposition 2.1] and it implies that $\Gamma$ is subgroup of $\mathrm{Aut}(A,\cdot)$ by the normality of $\mathrm{Inn}(A,\cdot)$.
Proposition 2.1.
We have the equalities
$\Gamma=\mathrm{Im}(\lambda)\mathrm{Inn}(A,\cdot)=\mathrm{Im}(\rho)\mathrm{Inn}(A,\cdot).$
Proof.
For any $a,b\in A$, a simple calculation using (2.1) yields that
$\lambda_{a}\cdot\rho_{b}$ $=\lambda_{a}\mathrm{conj}(b)\lambda_{b}=\lambda_{a}\lambda_{b}\cdot\mathrm{conj}(\lambda_{b}^{-1}(b)),$
$\lambda_{a}\cdot\mathrm{conj}(b)$ $=\mathrm{conj}(a^{-1})\rho_{a}\mathrm{conj}(b)=\rho_{a}\cdot\mathrm{conj}(\rho_{a}^{-1}(a^{-1})b),$
$\rho_{a}\cdot\mathrm{conj}(b)$ $=\mathrm{conj}(\rho_{a}(b)^{-1})^{-1}\rho_{a}=\lambda_{\rho_{a}(b)^{-1}}\cdot\rho_{\rho_{a}(b)^{-1}}^{-1}\rho_{a}.$
The equalities now follow.∎
Using the homomorphisms $\lambda$ and $\rho$, there are two left ideals
$J_{1}:=\mathrm{conj}^{-1}(\mathrm{Im}(\lambda)),\quad J_{2}:=\ker(\rho)$
that we can construct. They are implicit in [1, Proof of Theorem A].
Proposition 2.2.
Both of $J_{1}$ and $J_{2}$ are left ideals of $A$.
Proof.
Clearly $J_{1}$ is a subgroup of $(A,\cdot)$. For any $a\in A$ and $j\in J_{1}$, we may write $\mathrm{conj}(j)=\lambda_{b}$ for some $b\in A$. It follows that
$\mathrm{conj}(\lambda_{a}(j))=\lambda_{a}\mathrm{conj}(j)\lambda_{a}^{-1}=\lambda_{a}\lambda_{b}\lambda_{a}^{-1}\in\mathrm{Im}(\lambda),$
which yields $\lambda_{a}(j)\in J_{1}$. This shows that $J_{1}$ is a left ideal of $A$.
Clearly $J_{2}$ is a subgroup of $(A,\circ)$. For any $a\in A$ and $j\in J_{2}$, we have
$j\circ a=\rho_{j}(a)\cdot j=a\cdot j,$
and so $J_{2}$ is also a subgroup of $(A,\cdot)$. The above further mplies that
$a\circ j\circ\overline{a}=a\circ(\overline{a}\cdot j)=(a\circ\overline{a})\cdot a^{-1}\cdot(a\circ j)=\lambda_{a}(j).$
Since $J_{2}$ is in fact a normal subgroup of $(A,\circ)$, we obtain $\lambda_{a}(j)\in J_{2}$. This proves that $J_{2}$ is a left ideal of $A$. ∎
3. Proof of Theorem 1.4
In this section, let $A=(A,\cdot,\circ)$ be a finite skew brace with $(A,\cdot)\simeq T^{n}$, where $T$ is a non-abelian simple group and $n\in\mathbb{N}$. Also, let
$\Gamma:=\mathrm{Im}(\lambda)\mathrm{Im}(\rho),\quad J_{1}:=\mathrm{conj}^{-1}(\mathrm{Im}(\lambda)),\quad J_{2}:=\ker(\rho)$
be defined as in the previous section. By Proposition 2.1, we have
$\mathrm{Inn}(A,\cdot)\leq\Gamma\leq\mathrm{Aut}(A,\cdot).$ (3.1)
By Proposition 2.2, we know that $J_{1}$ and $J_{2}$ are left ideals of $A$.
3.1. The case when $n=1$
First, suppose that $A$ is almost trivial. Then the left ideals of $A$ are exactly the normal subgroups of $(A,\cdot)$. But $(A,\cdot)$ is non-abelian simple, so clearly $A$ is left-simple.
Conversely, suppose that $A$ is left-simple. Then $A$ is non-trivial, namely, $a\circ b\neq a\cdot b$ for some $a,b\in A$, for otherwise any subgroup of $(A,\cdot)$ would be a left ideal of $A$. We suppose for contradiction that $A$ is not almost trivial. Since $A$ is neither trivial nor almost trivial, we deduce from [3] that $(A,\circ)$ is not simple (the result of [3] is stated in terms of Hopf–Galois structures, but one can easily translate it into skew braces, using [9], for example). We then have $(A,\circ)\not\simeq(A,\cdot)$. By order consideration, it follows that
$\mathrm{Im}(\lambda)\not\supseteq\mathrm{Inn}(A,\cdot)\quad\mbox{and}\quad\mathrm{Im}(\rho)\not\supseteq\mathrm{Inn}(A,\cdot).$ (3.2)
Observe that $J_{1}\neq A$ because $\mathrm{Im}(\lambda)\not\supseteq\mathrm{Inn}(A,\cdot)$, and $J_{2}\neq A$ because $A$ is not almost trivial. Since $J_{1},J_{2}$ are left ideals of $A$ by Proposition 2.2, the left-simplicity of $A$ implies that $J_{1}=J_{2}=1$.
Note that $\Gamma$ is an almost simple group with socle $(A,\cdot)\simeq T$ by (3.1). We consider its factorization $\Gamma=\mathrm{Im}(\lambda)\mathrm{Im}(\rho)$. By (3.2), both of the factors are core-free, in the sense that they do not contain the socle. Moreover:
  1. $\bullet$
    Since $J_{1}=1$, the factor $\mathrm{Im}(\lambda)$ intersects trivially with $\mathrm{Inn}(A,\cdot)$ and so it embeds into $\mathrm{Out}(A,\cdot)$. Since $\mathrm{Out}(A,\cdot)$ is solvable by Schreier Conjecture (a consequence of the Classification of Finite Simple Groups), we deduce that $\mathrm{Im}(\lambda)$ is also solvable. Also $\mathrm{Im}(\lambda)\neq 1$ because $A$ is non-trivial.
  2. $\bullet$
    Since $J_{2}=1$, the factor $\mathrm{Im}(\rho)$ is isomorphic to $(A,\circ)$.
Fortunately, the core-free factorizations of finite almost simple groups with a solvable factor have been classified in [8, Theorem 1.1]. Put $H:=\mathrm{Im}(\lambda)$, which is solvable, and $K:=\mathrm{Im}(\rho)$. Identifying $(A,\cdot)$ with $T$, we have:
  1. (1)
    $|H|\neq 1$,  $|H\cap\mathrm{Inn}(T)|=1$;
  2. (2)
    $|H|$ divides $|\mathrm{Out}(T)|$;
  3. (3)
    $|K|=|T|$.
From Proposition 2.1, we also know that
$\frac{K}{K\cap\mathrm{Inn}(T)}\simeq\frac{K\mathrm{Inn}(T)}{\mathrm{Inn}(T)}=\frac{H\mathrm{Inn}(T)}{\mathrm{Inn}(T)}\simeq\frac{H}{H\cap\mathrm{Inn}(T)}.$
Together with (1) and (3), this yields the equality:
  1. (4)
    $|T|=|H||K\cap\mathrm{Inn}(T)|$.
We now use [8] to show that all possibilities of $(H,K)$ are impossible.
Suppose first that $K$ is also solvable. By [8, Proposition 4.1], for
$(H^{\prime},K^{\prime})=(H,K)\quad\mbox{or}\quad(H^{\prime},K^{\prime})=(K,H),$
the next two tables give all of the possibilities.
$T$ $|K^{\prime}\cap\mathrm{Inn}(T)|$ is divisible by $|K^{\prime}\cap\mathrm{Inn}(T)|$ divides
$\mathrm{PSL}_{2}(q)$ $q$ $\frac{q(q-1)}{(2,q-1)}$
$T$ $|\mathrm{Out}(T)|$ $|H^{\prime}|$ is divisible by $|K^{\prime}|$ is divisible by
$\mathrm{PSL}_{2}(7)$ 22 77 $|\mathrm{S}_{4}|$
$\mathrm{PSL}_{2}(11)$ 22 $11\cdot 5$ $|\mathrm{A}_{4}|$
$\mathrm{PSL}_{2}(23)$ 22 $23\cdot 11$ $|\mathrm{S}_{4}|$
$\mathrm{PSL}_{3}(3)$ 22 $13$ $3^{2}\cdot 2\cdot|\mathrm{S}_{4}|$
$\mathrm{PSL}_{3}(3)$ 22 $13\cdot 3$ $|\mathrm{A}\Gamma\mathrm{L}_{1}(9)|$
$\mathrm{PSL}_{3}(4)$ $12$ $7\cdot 3\cdot|\mathrm{S}_{3}|$ $2^{4}\cdot 3\cdot|\mathrm{D}_{10}|\cdot 2$
$\mathrm{PSL}_{3}(8)$ 66 $73\cdot 9$ $2^{3+6}\cdot 7^{2}\cdot 3$
$\mathrm{PSU}_{3}(8)$ $18$ $57\cdot 9$ $2^{3+6}\cdot 63\cdot 3$
$\mathrm{PSU}_{4}(2)$ 22 $2^{4}\cdot 5$ $3^{1+2}\cdot 2\cdot|\mathrm{A}_{4}|$
$\mathrm{PSU}_{4}(2)$ 22 $2^{4}\cdot|\mathrm{D}_{10}|$ $3^{1+2}\cdot 2\cdot|\mathrm{A}_{4}|$
$\mathrm{PSU}_{4}(2)$ 22 $2^{4}\cdot 5\cdot 4$ $3^{3}\cdot|\mathrm{S}_{3}|$
$\mathrm{M}_{11}$ 11 $11\cdot 5$ $|\mathrm{M}_{9}|\cdot 2$
In the second table, observe that
$|\mathrm{A}\Gamma\mathrm{L}_{1}(9)|=144,\quad|\mathrm{D}_{10}|=10,\quad|\mathrm{M}_{9}|=72.$
We see that neither $|H^{\prime}|$ nor $|K^{\prime}|$ divides $|\mathrm{Out}(T)|$, which contradicts (2).
In the first table, we must have $(H^{\prime},K^{\prime})=(H,K)$ by (1). Note that
$|T|$ $=|\mathrm{PSL}_{2}(q)|=\frac{q(q-1)(q+1)}{(2,q-1)},$
$|\mathrm{Out}(T)|$ $=|\mathrm{Out}(\mathrm{PSL}_{2}(q))|=(2,q-1)f,$
where $q=p^{f}$ and $p$ is a prime. We then see from (2) and the third column of the table that $|H||K\cap\mathrm{Inn}(T)|$ has to divide
$(2,q-1)f\cdot\frac{q(q-1)}{(2,q-1)}=q(q-1)f,$
which is strictly smaller than $|T|$ because
$f<\frac{2^{f}+1}{2}\leq\frac{q+1}{(2,q-1)}.$
This yields a contradiction to (4).
We have shown that $K$ must be insolvable. From here, we argue depending on the type of $T$. Let us remark that $T$ cannot be an exceptional group of Lie type by [8, Proposition 4.2].
I. $T$ is an alternating group
[…]
II. $T$ is a sporadic group
[…]
III. $T$ is a classical group of Lie type and is not an alternating group
The possibilities for $(H,K)$ are given in [8, Tables 1.1 and 1.2]. However, they only give an upper bound for $H\cap\mathrm{Inn}(T)$, which is not enough for us. For [8, Table 1.1], we shall further apply [2, Theorem 1], which provides a lower bound for $|H|$ and we can derive a contradiction to (2). For [8, Table 1.2], we shall use the description of $K\cap\mathrm{Inn}(T)$ and obtain a contradiction to (4). We summarize all of the possibilities in the two tables below.
$T$ $|H|$ is divisible by
(a) $\mathrm{PSL}_{n}(q)$, $n\geq 2$ $\frac{q^{n}-1}{q-1}$
(b) $\mathrm{PSL}_{4}(q)$ $\frac{q^{3}(q^{3}-1)}{(2,q-1)}$
(c) $\mathrm{PSp}_{2m}(q)$, $m\geq 2$, $q$ even $q^{m}(q^{m}-1)$
(d) $\mathrm{PSp}_{4}(q)$, $q$ even $q^{2}(q^{2}-1)$
(e) $\mathrm{PSp}_{4}(q)$, $q$ odd $q^{3}(q^{2}-1)$
(f) $\mathrm{PSU}_{2m}(q)$, $m\geq 2$ $\frac{q^{2m}(q^{2m}-1)}{q+1}$
(g) $\Omega_{2m+1}(q)$, $m\geq 3$, $q$ odd $\frac{1}{2}q^{m(m+1)/2}(q^{m}-1)$
(h) $\mathrm{P}\Omega^{+}_{2m}(q)$, $m\geq 5$ $\frac{q^{m}(q^{m}-1)}{(2,q-1)}$
(i) $\mathrm{P}\Omega_{8}^{+}(q)$ $\frac{q^{4}(q^{4}-1)}{(2,q-1)}$
$T$ $|T|$ $|\mathrm{Out}(T)|$ $|K\cap\mathrm{Inn}(T)|$ divides
$\mathrm{PSL}_{2}(11)$ $660$ 22 $|\mathrm{A}_{5}|$
$\mathrm{PSL}_{2}(16)$ $4080$ 44 $|\mathrm{A}_{5}|$
$\mathrm{PSL}_{2}(19)$ $3420$ 22 $|\mathrm{A}_{5}|$
$\mathrm{PSL}_{2}(29)$ $12180$ 22 $|\mathrm{A}_{5}|$
$\mathrm{PSL}_{2}(59)$ $102660$ 22 $|\mathrm{A}_{5}|$
$\mathrm{PSL}_{4}(3)$ $6065280$ 44 $3^{3}\cdot|\mathrm{PSL}_{3}(3)|$
$\mathrm{PSL}_{4}(3)$ $6065280$ 44 $4\cdot|\mathrm{PSL}_{2}(9)|\cdot 2$
$\mathrm{PSL}_{4}(4)$ $987033600$ 44 $5\cdot|\mathrm{PSL}_{2}(16)|\cdot 2$
$\mathrm{PSL}_{5}(2)$ $9999360$ 22 $2^{6}\cdot|\mathrm{S}_{3}|\cdot|\mathrm{PSL}_{3}(2)|$
$\mathrm{PSp}_{4}(3)$ $25920$ 22 $2^{4}\cdot|\mathrm{A}_{5}|$
$\mathrm{PSp}_{4}(3)$ $25920$ 22 $2^{4}\cdot|\mathrm{S}_{6}|$
$\mathrm{PSp}_{4}(5)$ $4680000$ 22 $|\mathrm{PSL}_{2}(5^{2})|\cdot 2$
$\mathrm{PSp}_{4}(7)$ $138297600$ 22 $|\mathrm{PSL}_{2}(7^{2})|\cdot 2$
$\mathrm{PSp}_{4}(11)$ $12860654400$ 22 $|\mathrm{PSL}_{2}(11^{2})|\cdot 2$
$\mathrm{PSp}_{4}(23)$ $20674026236160$ 22 $|\mathrm{PSL}_{2}(23^{2})|\cdot 2$
$\mathrm{Sp}_{6}(2)$ $1451520$ 11 $|\mathrm{S}_{8}|$
$\mathrm{PSp}_{6}(3)$ $4585351680$ 22 $|\mathrm{PSL}_{2}(27)|\cdot 3$
$\mathrm{PSU}_{3}(3)$ $6048$ 22 $|\mathrm{PSL}_{2}(7)|$
$\mathrm{PSU}_{3}(5)$ $126000$ 66 $|\mathrm{A}_{7}|$
$\mathrm{PSU}_{4}(3)$ $3265920$ 88 $|\mathrm{PSL}_{3}(4)|$
$\mathrm{PSU}_{4}(8)$ $34693789777920$ 66 $2^{12}\cdot|\mathrm{SL}_{2}(64)|\cdot 7$
$\Omega_{7}(3)$ $4585351680$ 22 $|\mathrm{G}_{2}(3)|$
$\Omega_{7}(3)$ $4585351680$ 22 $|\mathrm{Sp}_{6}(2)|$
$\Omega_{9}(3)$ $65784756654489600$ 22 $|\Omega_{8}^{-}(3)|\cdot 2$
$\Omega_{8}^{+}(2)$ $174182400$ 66 $|\mathrm{Sp}_{6}(2)|$
$\Omega_{8}^{+}(2)$ $174182400$ 66 $|\mathrm{A}_{9}|$
$\mathrm{P}\Omega_{8}^{+}(3)$ $4952179814400$ $24$ $|\Omega_{7}(3)|$
$\mathrm{P}\Omega_{8}^{+}(3)$ $4952179814400$ $24$ $|\Omega_{8}^{+}(2)|$
In the second table, let $\mathcal{K}$ be the number in the last column. We deduce from (2) that $|H||K\cap\mathrm{Inn}(T)|$ has to divide $|\mathrm{Out}(T)|\mathcal{K}$. Note that:
$|\mathrm{PSL}_{3}(3)|$ $5616$ $|\mathrm{PSL}_{2}(7)|$ $168$
$|\mathrm{PSL}_{2}(9)|$ $360$ $|\mathrm{PSL}_{3}(4)|$ $20160$
$|\mathrm{PSL}_{2}(16)|$ $4080$ $|\mathrm{SL}_{2}(64)|$ $262080$
$|\mathrm{PSL}_{3}(2)|$ $168$ $|\mathrm{G}_{2}(3)|$ $4245696$
$|\mathrm{PSL}_{2}(5^{2})|$ $7800$ $|\mathrm{Sp}_{6}(2)|$ $1451520$
$|\mathrm{PSL}_{2}(7^{2})|$ $58800$ $|\Omega_{8}^{-}(3)|$ $10151968619520$
$|\mathrm{PSL}_{2}(11^{2})|$ $885720$ $|\Omega_{7}(3)|$ $4585351680$
$|\mathrm{PSL}_{2}(23^{2})|$ $74017680$ $|\Omega_{8}^{+}(2)|$ $174182400$
$|\mathrm{PSL}_{2}(27)|$ $9828$
It is then straightforward to check that
$|\mathrm{Out}(T)|\mathcal{K}<|T|$
in every row of the second table. This yields a contradiction to (4).
In the first table, write $q=p^{f}$, where $p$ is a prime. As listed in [8, Table 2.1], for example, we know that:
$T$ $|\mathrm{Out}(T)|$
$\mathrm{PSL}_{2}(q)$ $(2,q-1)\cdot f$
$\mathrm{PSL}_{n}(q)$, $n\geq 3$ $(n,q-1)\cdot 2\cdot f$
$\mathrm{PSU}_{n}(q)$, $n\geq 3$ $(n,q+1)\cdot 2f$
$\mathrm{PSp}_{2m}(q)$, $(m,p)\neq(2,2)$ $(2,q-1)\cdot f$
$\mathrm{PSp}_{4}(q)$, $q$ even $2f$
$\Omega_{2m+1}(q)$, $m\geq 3$, $q$ odd $2\cdot f$
$\mathrm{P}\Omega_{2m}^{+}(q)$, $m\geq 5$, $q^{m}\not\equiv 1\pmod{4}$ $2\cdot(2,q-1)\cdot f$
$\mathrm{P}\Omega_{2m}^{+}(q)$, $m\geq 5$, $q^{m}\equiv 1\pmod{4}$ $8\cdot f$
$\mathrm{P}\Omega_{8}^{+}(q)$ $(2+(2,q-1))!\cdot f$
For an integer $z$, write $v_{p}(z)$ for the $p$-adic valuation of $z$. Using the trivial bound that $v_{p}(2)\leq 1$ when $2$ divides $|\mathrm{Out}(T)|$, we see that:
$v_{p}(|\mathrm{Out}(T)|)$ is at most $v_{p}(|H|)$ is at least
(b) $1+v_{p}(f)$ $3f$
(c) $1+v_{p}(f)$, $p=2$ mfmf, $m\geq 2$
(d) $1+v_{p}(f)$, $p=2$ $2f$
(e) $v_{p}(f)$, $p\geq 3$ $3f$
(f) $1+v_{p}(f)$ $2mf$, $m\geq 2$
(g) $v_{p}(f)$, $p\geq 3$ $\frac{m(m+1)}{2}f$, $m\geq 3$
(h) $3+v_{p}(f)$ mfmf, $m\geq 5$
(i) $3+v_{p}(f)$ $4f$
[…]
$v_{p}(|\mathrm{Out}(T)|)<v_{p}(|H|).$
But then $|H|$ cannot divide $|\mathrm{Out}(T)|$, and this contradicts (2).
Finally, we deal with case (a). If $n=2$, then
$|\mathrm{Out}(T)|\leq 2f<2^{f}+1\leq p^{f}+1=\frac{q^{2}-1}{q-1}\leq|H|,$
which contradicts (2). If $n\geq 3$ with $(fn,p)=(6,2)$, then
$\begin{cases}|\mathrm{Out}(T)|=2<\frac{2^{6}-1}{2-1}\leq|H|&\mbox{for }(f,n)=(1,6),\,q=2,\\ |\mathrm{Out}(T)|=12<\frac{4^{3}-1}{4-1}\leq|H|&\mbox{for }(f,n)=(2,3),\,q=4,\end{cases}$
which again contradicts (2). If $n\geq 3$ with $(fn,p)\neq(6,2)$, then we deduce from Zsigmondy’s theorem that $p^{fn}-1$ has a primitive prime divisor $\ell$. In other words, the prime $\ell$ is such that
$\ell\mid p^{fn}-1,\quad\ell\nmid p^{k}-1\mbox{ for all }1\leq k\leq fn-1.$ (3.3)
It follows that $\ell$ divides $\frac{q^{n}-1}{q-1}$ and hence $|H|$. By (2), we see that $\ell$ divides
$|\mathrm{Out}(T)|=(n,q-1)\cdot 2\cdot f.$
Since $\ell\nmid q-1$ and clearly $\ell\neq 2$, this means that $\ell\mid f$. But $p\equiv p^{\ell}\pmod{\ell}$ so we deduce from the first condition in (3.3) that
$p^{\frac{f}{\ell}n}-1\equiv p^{fn}-1\equiv 0\pmod{\ell}$.
This contradicts the second condition in (3.3).
We have now ruled out every possibility. Therefore, the assumption that $A$ is not almost trivial is false. This completes the proof of Theorem 1.4(a).
3.2. The case when $n\geq 2$.
Suppose that $A$ is left-simple. Note that
$\lambda_{a}\equiv\rho_{a}\hskip-5.69054pt\pmod{\mathrm{Inn}(A,\cdot)}$ (3.4)
for all $a\in A$ by (2.1). Observe that
$\mathrm{Im}(\lambda)\not\subseteq\mathrm{Inn}(A,\cdot),\mbox{ or equivalently }\mathrm{Im}(\rho)\not\subseteq\mathrm{Inn}(A,\cdot),$
for otherwise $T\times 1\times\cdots\times 1$ would be a proper left ideal of $A$. We see that $J_{1}\neq A$ by order consideration, and $J_{2}\neq A$. But $J_{1},J_{2}$ are left ideals of $A$ by Proposition 2.2, so the left-simplicity of $A$ implies that $J_{1}=J_{2}=1$.
  1. $\bullet$
    Since $J_{1}=1$, the image $\mathrm{Im}(\lambda)$ of $\lambda$ intersects trivially with $\mathrm{Inn}(A,\cdot)$, as claimed in the theorem.
  2. $\bullet$
    Since $J_{2}=1$, the circle group $(A,\circ)$ is isomorphic to the subgroup $\mathrm{Im}(\rho)$ of $\mathrm{Aut}(A,\cdot)\simeq\mathrm{Aut}(T)^{n}\rtimes S_{n}$. The projection $P$ of $\mathrm{Im}(\rho)$ onto $S_{n}$ is equal to that of $\mathrm{Im}(\lambda)$ by (3.4). Thus, for any orbit $O$ of the natural action of $P$ on $\{1,2,\dots,n\}$, the subgroup of $(A,\cdot)$ defined by
    $\prod_{i\in O}T_{i},$
    where $T_{i}$ denotes the $i$th copy of $T$ in $(A,\cdot)\simeq T^{n}$, is a left ideal of $A$. But $A$ is left-simple, so necessarily $O=\{1,2,\dots,n\}$. This proves that $P$ is a transitive subgroup of $S_{n}$.
This completes the proof of Theorem 1.4(b).
4. Connection with minimal Hopf–Galois structures
Let $L/K$ be a finite extension of fields. A Hopf–Galois structure on $L/K$ is a finite-dimensional cocommutative Hopf $K$-algebra $\mathcal{H}$ that acts on $L$ via a $K$-algebra homomorphism $\mathcal{H}\longrightarrow\mathrm{End}_{K}(L)$ such that the action respects the counit $\epsilon$, multiplication $m$, and comultiplication $\Delta$, in the sense that
$h(1)=\epsilon(h)(1),\quad h(xy)=m(\Delta(h)(x\otimes y))$
for all $h\in\mathcal{H},\,x,y\in L$, and the natural $K$-homomorphism
$L\otimes_{K}\mathcal{H}\longrightarrow\mathrm{End}_{K}(L)$
is bijective. The notion of a Hopf–Galois structure was introduced by [4] (or see [5] for detailed overview). For each Hopf $K$-subalgebra $\mathcal{H}^{\prime}$ of $\mathcal{H}$, let
$L^{\mathcal{H}^{\prime}}=\{x\in L\mid h^{\prime}(x)=\epsilon(h^{\prime})x\mbox{ for all }h^{\prime}\in\mathcal{H}^{\prime}\}$
denote its fixed field. This yields a Hopf–Galois correspondence
[…]
$\mathcal{H^{\prime}}$ $\longmapsto L^{\mathcal{H}^{\prime}}$
that is analogous to the Galois correspondence
$\{\mbox{subgroups of $G$}\}$ $\longrightarrow\{\mbox{intermediate fields of $L/K$}\};$
$H$ $\longmapsto L^{H}$
in Galois theory. Indeed, when $L/K$ is a Galois extension with Galois group $G$, the group ring $K[G]$ is naturally a Hopf–Galois structure on $L/K$, and the map $\Phi_{K[G]}$ reduces to the above Galois correspondence.
It is known by [4] that the map $\Phi_{\mathcal{H}}$ is always injective, but it need not be surjective in general. We can of course ask when $\Phi_{\mathcal{H}}$ is surjective. Here, we consider the other extreme and ask when $\Phi_{\mathcal{H}}$ is as far from being surjective as possible, i.e. when $\mathrm{Im}(\Phi_{\mathcal{H}})$ only contains $L$ and $K$. Since $\Phi_{\mathcal{H}}$ is injective, this is precisely the situation when $K$ and $\mathcal{H}$ are the only $K$-subalgebras of $\mathcal{H}$. The next definition is due to [6].
Definition 4.1.
A Hopf–Galois structure $\mathcal{H}$ on $L/K$ is said to be minimal if $\dim_{K}(\mathcal{H})\neq 1$ and $\mathcal{H}$ has no Hopf $K$-subalgebras other than $K$ and $\mathcal{H}$.
The theory of Hopf–Galois structures applies to all extensions. But when restricted to separable extensions, we have a very nice group-theoretic classification of Hopf–Galois structures thanks to [7]. In addition, when further restricted to Galois extensions, the classification may be translated into the language of skew braces. The next two theorems are due to [9].
Theorem 4.2.
Let $L/K$ be any finite Galois extension of fields with Galois group $G=(G,\circ)$. There is a bijective correspondence between:
  1. $(1)$
    the Hopf–Galois structures $\mathcal{H}$ on $L/K$;
  2. $(2)$
    the binary operations $\cdot$ on $G$ such that $(G,\cdot,\circ)$ is a skew brace.
Specifically, such an operation $\cdot$ is associated to the Hopf–Galois structure
$L[(G,\cdot)]^{(G,\circ)}=\left\{\sum_{\tau\in G}\ell_{\tau}\tau:L[(G,\cdot)]:\sigma(\ell_{\tau})=\ell_{\lambda_{\sigma}(\tau)}\mbox{ for all }\sigma,\tau\in G\right\}$
whose action on $L$ is defined by
$\left(\sum_{\tau\in G}\ell_{\tau}\tau\right)(x)=\sum_{\tau\in G}\ell_{\tau}\tau(x)$
for all $x\in L$, where $\lambda$ is the lambda map (1.1) of the skew brace $(G,\cdot,\circ)$.
Theorem 4.3.
Let $L/K$ be any finite Galois extension of fields with Galois group $G=(G,\circ)$. Let $\cdot$ be any binary operation on $G$ such that $(G,\cdot,\circ)$ is a skew brace and let $\mathcal{H}$ be the associated Hopf–Galois structure on $L/K$. For any subgroup $I$ of $G$, the following are equivalent:
  1. $(1)$
    the intermediate field $L^{I}$ lies in the image of $\Phi_{\mathcal{H}}$;
  2. $(2)$
    the subgroup $I$ is a left ideal of $(G,\cdot,\circ)$.
In particular, the following are equivalent:
  1. $(a)$
    the Hopf–Galois structure $\mathcal{H}$ is minimal;
  2. $(b)$
    the skew brace $(G,\cdot,\circ)$ is left-simple.
Therefore, in the language of Hopf–Galois theory, our Theorem 1.2, Corollary 1.3, and Theorem 1.4 give a partial classification of the minimal Hopf–Galois structures on finite Galois extensions of fields.
Acknowledgements
This work is supported by JSPS KAKENHI 24K16891.