### v2

# Skew braces with no proper left ideals

Abstract.
A skew brace $A=(A,\cdot,\circ)$ is said to be left-simple if $A\neq 1$ and it has no left ideal other than $1$ and $A$. The purpose of this paper is to give a partial classification of the finite left-simple skew braces. A result of Stefanello and Trappeniers implies that finite left-simple skew braces correspond precisely to minimal Hopf–Galois structures on finite Galois extensions of fields.
Key words and phrases:
skew brace, left ideal, minimal Hopf–Galois structure
2020 Mathematics Subject Classification:
Primary 20N99 16Y99, Secondary 12F10
1. Introduction
A skew brace is a set $A=(A,\cdot,\circ)$ equipped with two group operations $\cdot$ and $\circ$ such that the so-called brace relation
$a\circ(b\cdot c)=(a\circ b)\cdot a^{-1}\cdot(a\circ c)$
holds for all $a,b,c\in A$. For any $a\in A$, we write $a^{-1}$ and $\overline{a}$, respectively, for the inverses of $a$ in the groups $(A,\cdot)$ and $(A,\circ)$. We easily check that $(A,\cdot)$ and $(A,\circ)$ share the same identity $1$. For each $a\in A$, it is straightforward to verify that the map
$\lambda_{a}:A\longrightarrow A;\quad\lambda_{a}(x)=a^{-1}\cdot(a\circ x)$
is an automorphism of $(A,\cdot)$, and it is also not hard to show that
$\lambda:(A,\circ)\longrightarrow\mathrm{Aut}(A,\cdot);\quad a\mapsto\lambda_{a}$ (1.1)
is a group homomorphism. A left ideal of $A$ is a subgroup $I$ of $(A,\cdot)$ that is invariant under the action of $\lambda$, namely $\lambda_{a}(I)\subseteq I$ for all $a\in A$. Note that $I$ is a subgroup of $(A,\circ)$ in this case. We introduce a new definition.
Definition 1.1.
We shall say that a skew brace $A=(A,\cdot,\circ)$ is left-simple if $A\neq 1$ and it has no left ideal other than $1$ and $A$.
Analogous to the Classification of Finite Simple Groups, it is natural to ask for a classification of the finite left-simple skew braces. The purpose of this paper is to give a partial classification.
Let $A=(A,\cdot,\circ)$ be any finite left-simple skew brace. The additive group of $A$, namely $(A,\cdot)$, must be characteristically simple because its characteristic subgroups are all left ideals of $A$. Hence, we have $(A,\cdot)\simeq T^{n}$ for some finite simple group $T$ and $n\in\mathbb{N}$. We are able to prove a complete classification when $T$ is abelian and when $n=1$. We give a few fairly restrictive conditions that $A$ must satisfy when $T$ is non-abelian and $n\geq 2$.
Theorem 1.2.
Let $A=(A,\cdot,\circ)$ be a finite skew brace such that $(A,\cdot)\simeq C_{p}^{n}$ for some prime $p$ and $n\in\mathbb{N}$. Then $A$ is left-simple if and only if $n=1$.
Proof.
We may regard the group homomorphism
$\lambda:(A,\circ)\longrightarrow\mathrm{Aut}(A,\cdot)\simeq\mathrm{GL}_{n}(\mathbb{F}_{p})$
as a linear representation of the finite $p$-group $(A,\circ)$ over the field $\mathbb{F}_{p}$ that has characteristic $p$. The left ideals of the skew brace $A$ then coincide with the $\lambda$-invariant subspaces of the vector space $(A,\cdot)\simeq\mathbb{F}_{p}^{n}$. This means that $A$ is left-simple if and only if $\lambda$ is irreducible. But it is well-known in modular representation theory (e.g. see [13, Proposition 6.2.1]) that $\lambda$ is irreducible if and only if it is the trivial representation. This proves the theorem. ∎
The following corollary is immediate from Theorem 1.2.
Corollary 1.3.
Up to isomorphism, the finite left-simple skew braces with an abelian additive group are exactly $\mathbb{F}_{p}=(\mathbb{F}_{p},+,+)$, where $p$ is any prime and ++ is the usual addition.
Theorem 1.4.
Let $A=(A,\cdot,\circ)$ be a finite skew brace such that $(A,\cdot)\simeq T^{n}$ for some non-abelian simple group $T$ and $n\in\mathbb{N}$. It is well-known that
$\mathrm{Aut}(T^{n})=\mathrm{Aut}(T)\wr S_{n}=\mathrm{Aut}(T)^{n}\rtimes S_{n},$
where $S_{n}$ denotes the symmetric group on $n$ letters.
  1. $(a)$
    In the case $n=1$, we have $A$ is left-simple if and only if $A$ is an almost trivial skew brace, namely, $a\circ b=b\cdot a$ for all $a,b\in A$.
  2. $(b)$
    In the case $n\geq 2$, if $A$ is left-simple, then $(A,\circ)$ is isomorphic to some subgroup of $\mathrm{Aut}(T^{n})$ whose projection onto $S_{n}$ is a transitive subgroup, and $\mathrm{Im}(\lambda)$ intersects trivially with $\mathrm{Inn}(A,\cdot)$.
Remark 1.5.
The motivation of this paper came from the study of minimal Hopf–Galois structures on finite Galois extensions. In the language of Hopf–Galois theory, our results above yield a partial classification of the minimal Hopf–Galois structures on finite Galois extensions (see Section 4).
Remark 1.6.
A skew brace $A=(A,\cdot,\circ)$ is said to be simple if $A\neq 1$ and it has no ideal other than $1$ and $A$. Recall that an ideal of $A$ is a left ideal $I$ of $A$ that is normal in both $(A,\cdot)$ and $(A,\circ)$. In this case, we can naturally define a quotient skew brace structure on the set
$A/I=\{a\cdot I:a\in A\}=\{a\circ I:a\in A\}.$
Thus, ideals in a skew brace are a natural analog of normal subgroups in a group, and simple skew braces are arguably a closer analog of simple groups than left-simple skew braces. A classification of finite simple skew braces is desirable, but it seems to be out of reach at the moment. For example, in a simple skew brace $A=(A,\cdot,\circ)$, it is possible that $(A,\cdot)$ is abelian [1].
2. Preliminaries
Let $A=(A,\cdot,\circ)$ be a skew brace. For each $a\in A$, the maps
$\lambda_{a}:A\longrightarrow A;$ $\quad\lambda_{a}(x)=a^{-1}\cdot(a\circ x)$
$\rho_{a}:A\longrightarrow A;$ $\quad\rho_{a}(x)=(a\circ x)\cdot a^{-1}$
are automorphisms of $(A,\cdot)$. Also, let
$\mathrm{conj}(a):A\longrightarrow A;\quad\mathrm{conj}(a)(x)=a\cdot x\cdot a^{-1}$
denote the inner automorphism of $(A,\cdot)$ induced by $a$, so that
$\rho_{a}=\mathrm{conj}(a)\lambda_{a},\quad\lambda_{a}=\mathrm{conj}(a^{-1})\rho_{a},\quad\mathrm{conj}(a)=\rho_{a}\lambda_{a}^{-1}$ (2.1)
hold. It is well-known that
$\lambda:(A,\circ)\longrightarrow\mathrm{Aut}(A,\cdot);$ $\quad a\mapsto\lambda_{a}$
$\rho:(A,\circ)\longrightarrow\mathrm{Aut}(A,\cdot);$ $\quad a\mapsto\rho_{a}$
are group homomorphisms. Also let
$\mathrm{conj}:(A,\cdot)\longrightarrow\mathrm{Aut}(A,\cdot);\quad a\mapsto\mathrm{conj}(a)$
denote the natural group homomorphism.
For the homomorphisms $\lambda$ and $\rho$, we consider the product
$\Gamma:=\mathrm{Im}(\lambda)\mathrm{Im}(\rho)$
of their images. The next proposition is essentially [12, Proposition 2.1] and it implies that $\Gamma$ is subgroup of $\mathrm{Aut}(A,\cdot)$ by the normality of $\mathrm{Inn}(A,\cdot)$.
Proposition 2.1.
We have the equalities
$\Gamma=\mathrm{Im}(\lambda)\mathrm{Inn}(A,\cdot)=\mathrm{Im}(\rho)\mathrm{Inn}(A,\cdot).$
Proof.
For any $a,b\in A$, a simple calculation using (2.1) yields that
$\lambda_{a}\cdot\rho_{b}$ $=\lambda_{a}\mathrm{conj}(b)\lambda_{b}=\lambda_{a}\lambda_{b}\cdot\mathrm{conj}(\lambda_{b}^{-1}(b)),$
$\lambda_{a}\cdot\mathrm{conj}(b)$ $=\mathrm{conj}(a^{-1})\rho_{a}\mathrm{conj}(b)=\rho_{a}\cdot\mathrm{conj}(\rho_{a}^{-1}(a^{-1})b),$
$\rho_{a}\cdot\mathrm{conj}(b)$ $=\mathrm{conj}(\rho_{a}(b)^{-1})^{-1}\rho_{a}=\lambda_{\rho_{a}(b)^{-1}}\cdot\rho_{\rho_{a}(b)^{-1}}^{-1}\rho_{a}.$
The equalities now follow.∎
Using the homomorphisms $\lambda$ and $\rho$, there are two left ideals
$J_{1}:=\mathrm{conj}^{-1}(\mathrm{Im}(\lambda)),\quad J_{2}:=\ker(\rho)$
that we can construct. They are implicit in [2, Proof of Theorem A].
Proposition 2.2.
Both of $J_{1}$ and $J_{2}$ are left ideals of $A$.
Proof.
Clearly $J_{1}$ is a subgroup of $(A,\cdot)$. Let $j\in J_{1}$ and write $\mathrm{conj}(j)=\lambda_{b}$ for some $b\in A$. For any $a\in A$, it then follows that
$\mathrm{conj}(\lambda_{a}(j))=\lambda_{a}\mathrm{conj}(j)\lambda_{a}^{-1}=\lambda_{a}\lambda_{b}\lambda_{a}^{-1}\in\mathrm{Im}(\lambda),$
which yields $\lambda_{a}(j)\in J_{1}$. This shows that $J_{1}$ is a left ideal of $A$.
Clearly $J_{2}$ is a subgroup of $(A,\circ)$. For any $a\in A$ and $j\in J_{2}$, we have
$j\circ a=\rho_{j}(a)\cdot j=a\cdot j,$
and so $J_{2}$ is also a subgroup of $(A,\cdot)$. The above further implies that
$a\circ j\circ\overline{a}=a\circ(\overline{a}\cdot j)=(a\circ\overline{a})\cdot a^{-1}\cdot(a\circ j)=\lambda_{a}(j).$
Since $J_{2}$ is in fact a normal subgroup of $(A,\circ)$, we obtain $\lambda_{a}(j)\in J_{2}$. This proves that $J_{2}$ is a left ideal of $A$. ∎
3. Proof of Theorem 1.4
In this section, let $A=(A,\cdot,\circ)$ be a finite skew brace with $(A,\cdot)\simeq T^{n}$, where $T$ is a non-abelian simple group and $n\in\mathbb{N}$. Also, let
$\Gamma:=\mathrm{Im}(\lambda)\mathrm{Im}(\rho),\quad J_{1}:=\mathrm{conj}^{-1}(\mathrm{Im}(\lambda)),\quad J_{2}:=\ker(\rho)$
be defined as in the previous section. By Proposition 2.1, we have
$\mathrm{Inn}(A,\cdot)\leq\Gamma\leq\mathrm{Aut}(A,\cdot).$ (3.1)
By Proposition 2.2, we know that $J_{1}$ and $J_{2}$ are left ideals of $A$.
3.1. The case when $n=1$
First, suppose that $A$ is almost trivial. Then the left ideals of $A$ are exactly the normal subgroups of $(A,\cdot)$. But $(A,\cdot)$ is non-abelian simple, so clearly $A$ is left-simple.
Conversely, suppose that $A$ is left-simple. Then $A$ is non-trivial, namely, $a\circ b\neq a\cdot b$ for some $a,b\in A$, for otherwise any subgroup of $(A,\cdot)$ would be a left ideal of $A$. We suppose for contradiction that $A$ is not almost trivial. Since $A$ is neither trivial nor almost trivial, we obtain from [4, Theorem 1.1] that $(A,\circ)$ is not simple (the result of [4] is stated in terms of Hopf–Galois structures, but using [11, Theorem 3.1], for example, to translate it into the language of skew braces, it says that a finite skew brace with a non-abelian simple multiplicative group is either trivial or almost trivial). We then have $(A,\circ)\not\simeq(A,\cdot)$. By order consideration, it follows that
$\mathrm{Im}(\lambda)\not\supseteq\mathrm{Inn}(A,\cdot)\quad\mbox{and}\quad\mathrm{Im}(\rho)\not\supseteq\mathrm{Inn}(A,\cdot).$ (3.2)
Indeed, note that $\mathrm{Inn}(A,\cdot)\simeq(A,\cdot)$ has order $|A|$, while
$\mathrm{Im}(\lambda)\simeq(A,\circ)/\ker(\lambda)\quad\mbox{and}\quad\mathrm{Im}(\rho)\simeq(A,\circ)/\ker(\rho)$
have order at most $|A|$. This implies that
[…]
and we have a contradiction in both cases.
Observe that $J_{1}\neq A$ because $\mathrm{Im}(\lambda)\not\supseteq\mathrm{Inn}(A,\cdot)$, and $J_{2}\neq A$ because $A$ is not almost trivial. Since $J_{1},J_{2}$ are left ideals of $A$ by Proposition 2.2, the left-simplicity of $A$ implies that $J_{1}=J_{2}=1$.
Note that $\Gamma$ is an almost simple group with socle $(A,\cdot)\simeq T$ by (3.1). We consider its factorization $\Gamma=\mathrm{Im}(\lambda)\mathrm{Im}(\rho)$. By (3.2), both of the factors are core-free, in the sense that they do not contain the socle. Moreover:
  1. $\bullet$
    Since $J_{1}=1$, the factor $\mathrm{Im}(\lambda)$ intersects trivially with $\mathrm{Inn}(A,\cdot)$ and so it embeds into $\mathrm{Out}(A,\cdot)$. Since $\mathrm{Out}(A,\cdot)$ is solvable by Schreier Conjecture (a consequence of the Classification of Finite Simple Groups), we deduce that $\mathrm{Im}(\lambda)$ is also solvable. Also $\mathrm{Im}(\lambda)\neq 1$ because $A$ is non-trivial.
  2. $\bullet$
    Since $J_{2}=1$, the factor $\mathrm{Im}(\rho)$ is isomorphic to $(A,\circ)$.
Fortunately, the core-free factorizations of finite almost simple groups with a solvable factor have been classified in [9, Theorem 1.1]. Put $H:=\mathrm{Im}(\lambda)$, which is solvable, and $K:=\mathrm{Im}(\rho)$. Identifying $(A,\cdot)$ with $T$, we have:
  1. (1)
    $|H|\neq 1$,  $|H\cap\mathrm{Inn}(T)|=1$;
  2. (2)
    $|H|$ divides $|\mathrm{Out}(T)|$;
  3. (3)
    $|K|=|T|$.
From Proposition 2.1, we also know that
$\frac{K}{K\cap\mathrm{Inn}(T)}\simeq\frac{K\mathrm{Inn}(T)}{\mathrm{Inn}(T)}=\frac{H\mathrm{Inn}(T)}{\mathrm{Inn}(T)}\simeq\frac{H}{H\cap\mathrm{Inn}(T)}.$
Together with (1) and (3), this yields the equality:
  1. (4)
    $|T|=|H||K\cap\mathrm{Inn}(T)|$.
We now use [9] to show that all candidates of $(H,K)$ are impossible.
Remark 3.1.
In the following tables, the orders of certain groups are listed. For the linear groups, they can be constructed in Magma via
  1. $\bullet$
    PSL(n,q) for the projective special linear group on $\mathrm{GF}(q)^{n}$;
  2. $\bullet$
    PSU(n,q) for the projective special unitary group on $\mathrm{GF}(q^{2})^{n}$;
  3. $\bullet$
    PSp(2n,q) for the projective sympletic group on $\mathrm{GF}(q)^{2n}$;
  4. $\bullet$
    PSO(2n+1,q) for the projective special orthogonal group on $\mathrm{GF}(q)^{2n+1}$;
    (the symbol $\mathrm{P}\Omega_{2n+1}(q)$ denotes its derived subgroup);
  5. $\bullet$
    PSOPlus(2n,q) and PSOMinus(2n,q), respectively, for the projective special orthogonal groups on $\mathrm{GF}(q)^{2n}$ with Witt indices $n$ and n1n-1;
    (the symbols $\mathrm{P}\Omega_{2n}^{+}(q)$ and $\mathrm{P}\Omega_{2n}^{-}(q)$ denote their derived subgroups);
  6. $\bullet$
    AGammaL(n,q) for the affine semilinear group on $\mathrm{GF}(q)^{n}$;
  7. $\bullet$
    ChevalleyGroup("G",2,q) for the Chevalley group or exceptional group of type $G_{2}$ over the field $\mathrm{GF}(q)$.
For the non-projective linear groups, simply remove P in the above. For the Mathieu groups, they can be constructed in GAP via
  1. $\bullet$
    MathieuGroup(n) for the Mathieu group acting on $n$ objects.
The orders of these groups may then be computed. The orders of the outer automorphism groups $\mathrm{Out}(T)$ may be computed via
$|\mathrm{Out}(T)|=|\mathrm{Aut}(T)|/|\mathrm{Inn}(T)|=|\mathrm{Aut}(T)|/|T|.$
For the symmetric and alternating groups, their orders are obvious, so we do not list them explicitly.
Suppose first that $K$ is also solvable. By [9, Proposition 4.1], for
$(H^{\prime},K^{\prime})=(H,K)\quad\mbox{or}\quad(H^{\prime},K^{\prime})=(K,H),$
the next two tables give all of the possibilities.
$T$ $|K^{\prime}\cap\mathrm{Inn}(T)|$ is divisible by $|K^{\prime}\cap\mathrm{Inn}(T)|$ divides
$\mathrm{PSL}_{2}(q)$ $q$ $\frac{q(q-1)}{(2,q-1)}$
$T$ $|\mathrm{Out}(T)|$ $|H^{\prime}|$ is divisible by $|K^{\prime}|$ is divisible by
$\mathrm{PSL}_{2}(7)$ 22 77 $|\mathrm{S}_{4}|$
$\mathrm{PSL}_{2}(11)$ 22 $11\cdot 5$ $|\mathrm{A}_{4}|$
$\mathrm{PSL}_{2}(23)$ 22 $23\cdot 11$ $|\mathrm{S}_{4}|$
$\mathrm{PSL}_{3}(3)$ 22 $13$ $3^{2}\cdot 2\cdot|\mathrm{S}_{4}|$
$\mathrm{PSL}_{3}(3)$ 22 $13\cdot 3$ $|\mathrm{A}\Gamma\mathrm{L}_{1}(9)|$
$\mathrm{PSL}_{3}(4)$ $12$ $7\cdot 3\cdot|\mathrm{S}_{3}|$ $2^{4}\cdot 3\cdot|\mathrm{D}_{10}|\cdot 2$
$\mathrm{PSL}_{3}(8)$ 66 $73\cdot 9$ $2^{3+6}\cdot 7^{2}\cdot 3$
$\mathrm{PSU}_{3}(8)$ $18$ $57\cdot 9$ $2^{3+6}\cdot 63\cdot 3$
$\mathrm{PSU}_{4}(2)$ 22 $2^{4}\cdot 5$ $3^{1+2}\cdot 2\cdot|\mathrm{A}_{4}|$
$\mathrm{PSU}_{4}(2)$ 22 $2^{4}\cdot|\mathrm{D}_{10}|$ $3^{1+2}\cdot 2\cdot|\mathrm{A}_{4}|$
$\mathrm{PSU}_{4}(2)$ 22 $2^{4}\cdot 5\cdot 4$ $3^{3}\cdot|\mathrm{S}_{3}|$
$\mathrm{M}_{11}$ 11 $11\cdot 5$ $|\mathrm{M}_{9}|\cdot 2$
In the second table, observe that
$|\mathrm{A}\Gamma\mathrm{L}_{1}(9)|=144,\quad|\mathrm{D}_{10}|=10,\quad|\mathrm{M}_{9}|=72.$
We see that neither $|H^{\prime}|$ nor $|K^{\prime}|$ divides $|\mathrm{Out}(T)|$, which contradicts (2).
In the first table, we must have $(H^{\prime},K^{\prime})=(H,K)$ by (1). Note that
$|T|$ $=|\mathrm{PSL}_{2}(q)|=\frac{q(q-1)(q+1)}{(2,q-1)},$
$|\mathrm{Out}(T)|$ $=|\mathrm{Out}(\mathrm{PSL}_{2}(q))|=(2,q-1)f,$
where $q=p^{f}$ and $p$ is a prime. We then see from (2) and the third column of the table that $|H||K\cap\mathrm{Inn}(T)|$ has to divide
$(2,q-1)f\cdot\frac{q(q-1)}{(2,q-1)}=q(q-1)f,$
which is strictly smaller than $|T|$ because
$f<\frac{2^{f}+1}{2}\leq\frac{q+1}{(2,q-1)}.$
This yields a contradiction to (4).
We have shown that $K$ must be insolvable. From here, we argue depending on the type of $T$. Let us remark that $T$ cannot be an exceptional group of Lie type by [9, Proposition 4.2].
I. $T$ is an alternating group
[…]
II. $T$ is a sporadic group
[…]
III. $T$ is a classical group of Lie type and is not an alternating group
The possibilities for $(H,K)$ are given in [9, Tables 1.1 and 1.2]. However, they only give an upper bound for $H\cap\mathrm{Inn}(T)$, which is not enough for us. For [9, Table 1.1], we shall further apply [3, Theorem 1], which provides a lower bound for $|H|$ and we can derive a contradiction to (2). For [9, Table 1.2], we shall use the description of $K\cap\mathrm{Inn}(T)$ and obtain a contradiction to (4). We summarize all of the possibilities in the two tables below.
$T$ $|H|$ is divisible by
(a) $\mathrm{PSL}_{n}(q)$, $n\geq 2$ $\frac{q^{n}-1}{q-1}$
(b) $\mathrm{PSL}_{4}(q)$ $\frac{q^{3}(q^{3}-1)}{(2,q-1)}$
(c) $\mathrm{PSp}_{2m}(q)$, $m\geq 2$, $q$ even $q^{m}(q^{m}-1)$
(d) $\mathrm{PSp}_{4}(q)$, $q$ even $q^{2}(q^{2}-1)$
(e) $\mathrm{PSp}_{4}(q)$, $q$ odd $q^{3}(q^{2}-1)$
(f) $\mathrm{PSU}_{2m}(q)$, $m\geq 2$ $\frac{q^{2m}(q^{2m}-1)}{q+1}$
(g) $\Omega_{2m+1}(q)$, $m\geq 3$, $q$ odd $\frac{1}{2}q^{m(m+1)/2}(q^{m}-1)$
(h) $\mathrm{P}\Omega^{+}_{2m}(q)$, $m\geq 5$ $\frac{q^{m}(q^{m}-1)}{(2,q-1)}$
(i) $\mathrm{P}\Omega_{8}^{+}(q)$ $\frac{q^{4}(q^{4}-1)}{(2,q-1)}$
$T$ $|T|$ $|\mathrm{Out}(T)|$ $|K\cap\mathrm{Inn}(T)|$ divides
$\mathrm{PSL}_{2}(11)$ $660$ 22 $|\mathrm{A}_{5}|$
$\mathrm{PSL}_{2}(16)$ $4080$ 44 $|\mathrm{A}_{5}|$
$\mathrm{PSL}_{2}(19)$ $3420$ 22 $|\mathrm{A}_{5}|$
$\mathrm{PSL}_{2}(29)$ $12180$ 22 $|\mathrm{A}_{5}|$
$\mathrm{PSL}_{2}(59)$ $102660$ 22 $|\mathrm{A}_{5}|$
$\mathrm{PSL}_{4}(3)$ $6065280$ 44 $3^{3}\cdot|\mathrm{PSL}_{3}(3)|$
$\mathrm{PSL}_{4}(3)$ $6065280$ 44 $4\cdot|\mathrm{PSL}_{2}(9)|\cdot 2$
$\mathrm{PSL}_{4}(4)$ $987033600$ 44 $5\cdot|\mathrm{PSL}_{2}(16)|\cdot 2$
$\mathrm{PSL}_{5}(2)$ $9999360$ 22 $2^{6}\cdot|\mathrm{S}_{3}|\cdot|\mathrm{PSL}_{3}(2)|$
$\mathrm{PSp}_{4}(3)$ $25920$ 22 $2^{4}\cdot|\mathrm{A}_{5}|$
$\mathrm{PSp}_{4}(3)$ $25920$ 22 $2^{4}\cdot|\mathrm{S}_{6}|$
$\mathrm{PSp}_{4}(5)$ $4680000$ 22 $|\mathrm{PSL}_{2}(5^{2})|\cdot 2$
$\mathrm{PSp}_{4}(7)$ $138297600$ 22 $|\mathrm{PSL}_{2}(7^{2})|\cdot 2$
$\mathrm{PSp}_{4}(11)$ $12860654400$ 22 $|\mathrm{PSL}_{2}(11^{2})|\cdot 2$
$\mathrm{PSp}_{4}(23)$ $20674026236160$ 22 $|\mathrm{PSL}_{2}(23^{2})|\cdot 2$
$\mathrm{Sp}_{6}(2)$ $1451520$ 11 $|\mathrm{S}_{8}|$
$\mathrm{PSp}_{6}(3)$ $4585351680$ 22 $|\mathrm{PSL}_{2}(27)|\cdot 3$
$\mathrm{PSU}_{3}(3)$ $6048$ 22 $|\mathrm{PSL}_{2}(7)|$
$\mathrm{PSU}_{3}(5)$ $126000$ 66 $|\mathrm{A}_{7}|$
$\mathrm{PSU}_{4}(3)$ $3265920$ 88 $|\mathrm{PSL}_{3}(4)|$
$\mathrm{PSU}_{4}(8)$ $34693789777920$ 66 $2^{12}\cdot|\mathrm{SL}_{2}(64)|\cdot 7$
$\Omega_{7}(3)$ $4585351680$ 22 $|\mathrm{G}_{2}(3)|$
$\Omega_{7}(3)$ $4585351680$ 22 $|\mathrm{Sp}_{6}(2)|$
$\Omega_{9}(3)$ $65784756654489600$ 22 $|\Omega_{8}^{-}(3)|\cdot 2$
$\Omega_{8}^{+}(2)$ $174182400$ 66 $|\mathrm{Sp}_{6}(2)|$
$\Omega_{8}^{+}(2)$ $174182400$ 66 $|\mathrm{A}_{9}|$
$\mathrm{P}\Omega_{8}^{+}(3)$ $4952179814400$ $24$ $|\Omega_{7}(3)|$
$\mathrm{P}\Omega_{8}^{+}(3)$ $4952179814400$ $24$ $|\Omega_{8}^{+}(2)|$
In the second table, let $\mathcal{K}$ be the number in the last column. We deduce from (2) that $|H||K\cap\mathrm{Inn}(T)|$ has to divide $|\mathrm{Out}(T)|\mathcal{K}$. Note that:
$|\mathrm{PSL}_{3}(3)|$ $5616$ $|\mathrm{PSL}_{2}(7)|$ $168$
$|\mathrm{PSL}_{2}(9)|$ $360$ $|\mathrm{PSL}_{3}(4)|$ $20160$
$|\mathrm{PSL}_{2}(16)|$ $4080$ $|\mathrm{SL}_{2}(64)|$ $262080$
$|\mathrm{PSL}_{3}(2)|$ $168$ $|\mathrm{G}_{2}(3)|$ $4245696$
$|\mathrm{PSL}_{2}(5^{2})|$ $7800$ $|\mathrm{Sp}_{6}(2)|$ $1451520$
$|\mathrm{PSL}_{2}(7^{2})|$ $58800$ $|\Omega_{8}^{-}(3)|$ $10151968619520$
$|\mathrm{PSL}_{2}(11^{2})|$ $885720$ $|\Omega_{7}(3)|$ $4585351680$
$|\mathrm{PSL}_{2}(23^{2})|$ $74017680$ $|\Omega_{8}^{+}(2)|$ $174182400$
$|\mathrm{PSL}_{2}(27)|$ $9828$
It is then straightforward to check that
$|\mathrm{Out}(T)|\mathcal{K}<|T|$
in every row of the second table. This yields a contradiction to (4).
In the first table, write $q=p^{f}$, where $p$ is a prime. As listed in [9, Table 2.1], for example, we know that:
$T$ $|\mathrm{Out}(T)|$
$\mathrm{PSL}_{2}(q)$ $(2,q-1)\cdot f$
$\mathrm{PSL}_{n}(q)$, $n\geq 3$ $(n,q-1)\cdot 2\cdot f$
$\mathrm{PSU}_{n}(q)$, $n\geq 3$ $(n,q+1)\cdot 2f$
$\mathrm{PSp}_{2m}(q)$, $(m,p)\neq(2,2)$ $(2,q-1)\cdot f$
$\mathrm{PSp}_{4}(q)$, $q$ even $2f$
$\Omega_{2m+1}(q)$, $m\geq 3$, $q$ odd $2\cdot f$
$\mathrm{P}\Omega_{2m}^{+}(q)$, $m\geq 5$, $q^{m}\not\equiv 1\pmod{4}$ $2\cdot(2,q-1)\cdot f$
$\mathrm{P}\Omega_{2m}^{+}(q)$, $m\geq 5$, $q^{m}\equiv 1\pmod{4}$ $8\cdot f$
$\mathrm{P}\Omega_{8}^{+}(q)$ $(2+(2,q-1))!\cdot f$
For an integer $z$, write $v_{p}(z)$ for the $p$-adic valuation of $z$. Using the trivial bound that $v_{p}(2)\leq 1$ when $2$ divides $|\mathrm{Out}(T)|$, we see that:
$v_{p}(|\mathrm{Out}(T)|)$ is at most $v_{p}(|H|)$ is at least
(b) $1+v_{p}(f)$ $3f$
(c) $1+v_{p}(f)$, $p=2$ mfmf, $m\geq 2$
(d) $1+v_{p}(f)$, $p=2$ $2f$
(e) $v_{p}(f)$, $p\geq 3$ $3f$
(f) $1+v_{p}(f)$ $2mf$, $m\geq 2$
(g) $v_{p}(f)$, $p\geq 3$ $\frac{m(m+1)}{2}f$, $m\geq 3$
(h) $3+v_{p}(f)$ mfmf, $m\geq 5$
(i) $3+v_{p}(f)$ $4f$
[…]
$v_{p}(|\mathrm{Out}(T)|)<v_{p}(|H|).$
But then $|H|$ cannot divide $|\mathrm{Out}(T)|$, and this contradicts (2).
Finally, we deal with case (a). If $n=2$, then
$|\mathrm{Out}(T)|\leq 2f<2^{f}+1\leq p^{f}+1=\frac{q^{2}-1}{q-1}\leq|H|,$
which contradicts (2). If $n\geq 3$ with $(fn,p)=(6,2)$, then
$\begin{cases}|\mathrm{Out}(T)|=2<\frac{2^{6}-1}{2-1}\leq|H|&\mbox{for }(f,n)=(1,6),\,q=2,\\ |\mathrm{Out}(T)|=12<\frac{4^{3}-1}{4-1}\leq|H|&\mbox{for }(f,n)=(2,3),\,q=4,\end{cases}$
which again contradicts (2). If $n\geq 3$ with $(fn,p)\neq(6,2)$, then we deduce from Zsigmondy’s theorem [14] (also see [10, Theorem 3]) that $p^{fn}-1$ has a primitive prime divisor $\ell$. In other words, the prime $\ell$ is such that
$\ell\mid p^{fn}-1,\quad\ell\nmid p^{k}-1\mbox{ for all }1\leq k\leq fn-1.$ (3.3)
It follows that $\ell$ divides $\frac{q^{n}-1}{q-1}$ and hence $|H|$. By (2), we see that $\ell$ divides
$|\mathrm{Out}(T)|=(n,q-1)\cdot 2\cdot f.$
Since $\ell\nmid q-1$ and clearly $\ell\neq 2$, this means that $\ell\mid f$. But $p\equiv p^{\ell}\pmod{\ell}$ so we deduce from the first condition in (3.3) that
$p^{\frac{f}{\ell}n}-1\equiv p^{fn}-1\equiv 0\pmod{\ell}$.
This contradicts the second condition in (3.3).
We have now ruled out every possibility. Therefore, the assumption that $A$ is not almost trivial is false. This completes the proof of Theorem 1.4(a).
3.2. The case when $n\geq 2$.
Suppose that $A$ is left-simple. Note that
$\lambda_{a}\equiv\rho_{a}\hskip-5.69054pt\pmod{\mathrm{Inn}(A,\cdot)}$ (3.4)
for all $a\in A$ by (2.1), and that
$\mathrm{Im}(\lambda)\not\subseteq\mathrm{Inn}(A,\cdot),$ (3.5)
for otherwise $T\times 1\times\cdots\times 1$ would be a non-trivial proper left ideal of $A$. We then see that $J_{1}\neq A$ by order consideration. Indeed, if $J_{1}=A$, namely the inclusion $\mathrm{Inn}(A,\cdot)\subseteq\mathrm{Im}(\lambda)$ holds, then we must in fact have an equality because $\mathrm{Inn}(A,\cdot)\simeq(A,\cdot)$ has order $|A|$, while
$\mathrm{Im}(\lambda)\simeq(A,\circ)/\ker(\lambda)$
has order at most $|A|$. But this contradicts (3.5). From (3.4) and (3.5), we also obtain $J_{2}\neq A$. Since $J_{1},J_{2}$ are left ideals of $A$ by Proposition 2.2, the left-simplicity of $A$ implies that $J_{1}=J_{2}=1$.
Since $J_{1}=1$, the image $\mathrm{Im}(\lambda)$ of $\lambda$ intersects trivially with $\mathrm{Inn}(A,\cdot)$, as claimed in the theorem.
Since $J_{2}=1$, the multiplicative group $(A,\circ)$ embeds into
$\mathrm{Aut}(A,\cdot)\simeq\mathrm{Aut}(T)^{n}\rtimes S_{n}$
via $\rho$, and let $P$ denote the projection of $\mathrm{Im}(\rho)$ onto $S_{n}$. Since
$\mathrm{Inn}(A,\cdot)\simeq\mathrm{Inn}(T)^{n}\subseteq\mathrm{Aut}(T)^{n},$
the projection of $\mathrm{Im}(\lambda)$ onto $S_{n}$ is also equal to $P$ by (3.4). For any orbit $O$ of the natural action of $P$ on $\{1,2,\dots,n\}$, the subgroup of $(A,\cdot)$ given by
$\prod_{i\in O}T_{i},$
where $T_{i}$ denotes the $i$th copy of $T$ in $(A,\cdot)\simeq T^{n}$, is then a left ideal of $A$. But $A$ is left-simple, so necessarily $O=\{1,2,\dots,n\}$. This proves that $P$ is a transitive subgroup of $S_{n}$.
This completes the proof of Theorem 1.4(b).
Remark 3.2.
The condition that $(A,\circ)$ is isomorphic to a subgroup, $X$ say, of $\mathrm{Aut}(T^{n})$ whose projection onto $S_{n}$ is transitive is very restrictive. To see why, consider the natural homomorphism
$\phi:X\lhook\joinrel\xrightarrow{\quad}\mathrm{Aut}(T)^{n}\rtimes S_{n}\longrightarrow\mathrm{Out}(T)^{n}\rtimes S_{n}.$
Its kernel $\ker(\phi)$ lies in $\mathrm{Inn}(T)^{n}\simeq T^{n}$, and its image $\mathrm{Im}(\varphi)\simeq X/\ker(\phi)$ has the same projection onto $S_{n}$ as $X$, which is transitive. It follows that
$\exists R\leq T^{n}\mbox{ s.t. }n\mid[T^{n}:R]\mbox{\,\ and\,\ }[T^{n}:R]\mid|\mathrm{Out}(T)|^{n}\cdot n!.$ (3.6)
In particular, the prime factors of $n$ must all divide $|T|$.
For example, take $T=(T,\cdot)=(A_{5},\cdot)$. Note that
$|A_{5}|=60=2^{2}\cdot 3\cdot 5\quad\mbox{and}\quad|\mathrm{Out}(A_{5})|=2.$
We see right away that $(A_{5}^{n},\cdot,\circ)$ can be left-simple only when $n=2^{x}3^{y}5^{z}$, where $x,y,z\in\mathbb{N}_{\geq 0}$. Moreover, a skew brace $(A_{5}^{2},\cdot,\circ)$ cannot be left-simple because $A_{5}^{2}$ has no subgroup of index $2,4$, or $8$.
4. Connection with minimal Hopf–Galois structures
Let $L/K$ be a finite extension of fields. A Hopf–Galois structure on $L/K$ is a finite-dimensional cocommutative Hopf $K$-algebra $\mathcal{H}$ that acts on $L$ via a $K$-algebra homomorphism $\mathcal{H}\longrightarrow\mathrm{End}_{K}(L)$ such that the action respects the counit $\epsilon$, multiplication $m$, and comultiplication $\Delta$, in the sense that
$h(1)=\epsilon(h)(1),\quad h(xy)=m(\Delta(h)(x\otimes y))$
for all $h\in\mathcal{H},\,x,y\in L$, and the natural $K$-homomorphism
$L\otimes_{K}\mathcal{H}\longrightarrow\mathrm{End}_{K}(L)$
is bijective. The notion of a Hopf–Galois structure was introduced by [5] (or see [6] for detailed overview). For each Hopf $K$-subalgebra $\mathcal{H}^{\prime}$ of $\mathcal{H}$, let
$L^{\mathcal{H}^{\prime}}=\{x\in L:h^{\prime}(x)=\epsilon(h^{\prime})x\mbox{ for all }h^{\prime}\in\mathcal{H}^{\prime}\}$
denote its fixed field. This yields a Hopf–Galois correspondence
[…]
$\mathcal{H^{\prime}}$ $\longmapsto L^{\mathcal{H}^{\prime}}$
that is analogous to the Galois correspondence in Galois theory. Indeed, in the case that $L/K$ is a $G$-Galois extension, the group ring $K[G]$ is naturally a Hopf–Galois structure on $L/K$, and the map $\Phi_{K[G]}$ above reduces to the usual Galois correspondence
$\{\mbox{subgroups of $G$}\}$ $\longrightarrow\{\mbox{intermediate fields of $L/K$}\};$
$G^{\prime}$ $\longmapsto L^{G^{\prime}}$
because the Hopf $K$-subalgebras of $K[G]$ are exactly the group rings $K[G^{\prime}]$, where $G^{\prime}$ ranges over the subgroups of $G$.
It is known by [5] that the map $\Phi_{\mathcal{H}}$ is always injective, but it need not be surjective in general. We can of course ask when $\Phi_{\mathcal{H}}$ is surjective. Here, we consider the other extreme and ask when $\Phi_{\mathcal{H}}$ is as far from being surjective as possible, i.e. when $\mathrm{Im}(\Phi_{\mathcal{H}})$ only contains $L$ and $K$. Since $\Phi_{\mathcal{H}}$ is injective, this is precisely the case when $K$ and $\mathcal{H}$ are the only Hopf $K$-subalgebras of $\mathcal{H}$. The next definition is due to [7].
Definition 4.1.
A Hopf–Galois structure $\mathcal{H}$ on $L/K$ is said to be minimal if $\dim_{K}(\mathcal{H})\neq 1$ and $\mathcal{H}$ has no Hopf $K$-subalgebras other than $K$ and $\mathcal{H}$.
The theory of Hopf–Galois structures applies to all extensions. But when restricted to separable extensions, we have a very nice group-theoretic classification of Hopf–Galois structures thanks to [8]. In addition, when further restricted to Galois extensions, the classification may be translated into the
language of skew braces. The next theorem is from [11, Theorem 3.1].
Theorem 4.2.
Let $L/K$ be any finite Galois extension of fields with Galois group $G=(G,\circ)$. There is a bijective correspondence between:
  1. $(1)$
    the Hopf–Galois structures $\mathcal{H}$ on $L/K$;
  2. $(2)$
    the binary operations $\cdot$ on $G$ such that $(G,\cdot,\circ)$ is a skew brace.
Specifically, such an operation $\cdot$ is associated to the Hopf–Galois structure
$L[(G,\cdot)]^{(G,\circ)}=\left\{\sum_{\tau\in G}\ell_{\tau}\tau\in L[(G,\cdot)]:\sigma(\ell_{\tau})=\ell_{\lambda_{\sigma}(\tau)}\mbox{ for all }\sigma,\tau\in G\right\}$
whose action on $L$ is defined by
$\left(\sum_{\tau\in G}\ell_{\tau}\tau\right)(x)=\sum_{\tau\in G}\ell_{\tau}\tau(x)$
for all $x\in L$, where $\lambda$ is the lambda map (1.1) of the skew brace $(G,\cdot,\circ)$.
Let us remark that the Hopf–Galois structure corresponding to the trivial skew brace is the group ring $K[G]$ and is called the classical structure, while the Hopf–Galois structure corresponding to the almost trivial skew brace is called the canonical non-classical structure.
The next theorem is from [11, Corollary 4.1].
Theorem 4.3.
Let $L/K$ be any finite Galois extension of fields with Galois group $G=(G,\circ)$. Let $\cdot$ be any binary operation on $G$ such that $(G,\cdot,\circ)$ is a skew brace and let $\mathcal{H}$ be the associated Hopf–Galois structure on $L/K$. For any subgroup $I$ of $G$, the following are equivalent:
  1. $(1)$
    the intermediate field $L^{I}$ lies in the image of $\Phi_{\mathcal{H}}$;
  2. $(2)$
    the subgroup $I$ is a left ideal of $(G,\cdot,\circ)$.
In particular, the following are equivalent:
  1. $(a)$
    the Hopf–Galois structure $\mathcal{H}$ is minimal;
  2. $(b)$
    the skew brace $(G,\cdot,\circ)$ is left-simple.
Therefore, in the language of Hopf–Galois theory, our Theorem 1.2, Corollary 1.3, and Theorem 1.4 give a partial classification of the minimal Hopf–Galois structures on finite Galois extensions of fields.
Acknowledgements
This work is supported by JSPS KAKENHI 24K16891. The author would like to thank the referees for their valuable suggestions.